Sunset Sales has 6.8 percent coupon bonds on the market with 11 years left to maturity. The bonds make semiannual payments and currently sell for 98.6 percent of par ($986). What is the YTM
Answer #1

Number of periods = 11 * 2 = 22

Coupon = (6.8% of 1000) / 2 = 34

YTM = 6.98%

Keys to use in a financial calculator:

2nd P/Y 2

FV 1000

PV -986

N 22

PMT 34

CPT I/Y